OpenStudy (jazzyfa30):

help please f (theta) =sin(theta) and g (theta)= cos(theta).find the exact value of each function below if theta =60 degrees a)f(theta) b) g (theta/2) c f(2 theata) d g(2 theta) e) {f(theta)}^2 f) (-theta) g)(-theta)

OpenStudy (jazzyfa30):

@mathmale

OpenStudy (mathmale):

Where is the function that you're supposed to evaluate?

OpenStudy (jazzyfa30):

sorry forgot to post them

OpenStudy (mathmale):

Bet you can do this one: a) f(theta) Give it a try, please. Hint: How is f(theta) defined? Hint #2: what is the value of theta?

OpenStudy (jazzyfa30):

the value of theta is 60 so f(theta) is f(60)?????

OpenStudy (mathmale):

Yes, and how is f(theta) defined? Take a look at the problem statement.

OpenStudy (jazzyfa30):

sin (60) so i use my trig table???

OpenStudy (mathmale):

Yes, indeed.

OpenStudy (jazzyfa30):

which is squareroot 3/2

OpenStudy (mathmale):

Yes. Fine. I will not be able to stay on OpenStudy long enough to guide you through all of those problems. But for now, pick one or two of the harder ones and we'll work on them.

OpenStudy (jazzyfa30):

ok can we do e

OpenStudy (mathmale):

e) {f(theta)}^2 1. What is the value of f(theta)?

OpenStudy (jazzyfa30):

squareroot 3/2

OpenStudy (mathmale):

Hint: Please write that as (Sqrt(3) /2 or as Sqrt(3) ------ 2 for greater clarity. Fine. That is f(theta). Now what's the square of f(theta)? It's the same as the square of \[\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (jazzyfa30):

3/2

OpenStudy (mathmale):

Check your denominator.

OpenStudy (jazzyfa30):

3/4

OpenStudy (mathmale):

\[(\frac{ \sqrt{3} }{ 2 })^2=?\]

OpenStudy (mathmale):

that's better. One more problem?

OpenStudy (jazzyfa30):

what?:????

OpenStudy (mathmale):

I was offering to help you with one more problem, if you wish.

OpenStudy (jazzyfa30):

ohh lol c

OpenStudy (mathmale):

(c) f(2 theata) How is f(theta) defined? Write it out.

OpenStudy (jazzyfa30):

squareroot 3/2

OpenStudy (mathmale):

No...I was asking for the DEFINITION of f(theta), not the value of f at some angle.

OpenStudy (jazzyfa30):

ohhh 60

OpenStudy (mathmale):

Hint: g(theta) is defined as cos theta. f(theta) is defined as ... ?

OpenStudy (jazzyfa30):

sin theta

OpenStudy (mathmale):

Yes, and so sin (1*theta) = what, when theta = 60 degrees?

OpenStudy (mathmale):

sin (1*theta) = sin (1*60 deg) = sin (60 deg) = ?

OpenStudy (jazzyfa30):

squareroot 3/2

OpenStudy (mathmale):

Right. Now, what about sin (2*theta) = sin (2*60 deg) = sin ( ? ) = ?

OpenStudy (jazzyfa30):

120

OpenStudy (mathmale):

True, 2*60 deg = 120 deg, but I've asked you for sin (2*60 deg) = sin ( ? ) = ??

OpenStudy (jazzyfa30):

i have to go thanks

OpenStudy (mathmale):

You're welcome. Plesae come bak and finish this problem; that's important.

OpenStudy (jazzyfa30):

@TwoPointInfinity